OpenStudy (crashonce):

The amount of $P invested for 5 years at 4%p.a. compound interest, payable monthly is?

OpenStudy (danjs):

Do you remember the formula for the compiund interest ?

OpenStudy (crashonce):

no

OpenStudy (crashonce):

oh I=p(1+r)^n

OpenStudy (danjs):

Interest = principle(1-rate)^number periods that it?

OpenStudy (danjs):

monthly at 12 months in a year for 5 years, means 12*5 = 60 periods

OpenStudy (crashonce):

so how does it link with the question, im still not sure

OpenStudy (danjs):

PV*(1-r)^n = FV PV = Present Value FV = Future Value r = annual interest rate n = number of periods

OpenStudy (danjs):

Here, r = 0.04 n = 60 periods PV = $P

OpenStudy (danjs):

so you have $P * (1-0.04)^60 = Future Value

OpenStudy (crashonce):

um isn't it 1+0.04

OpenStudy (danjs):

sorry, you are right, my mistake $P * (1 + 0.04)^20 = Future Value

OpenStudy (danjs):

Is there more to the question?

OpenStudy (crashonce):

no thanks

OpenStudy (danjs):

welcome